OpenStudy (anonymous):

HELPP!! LET: \[P(X)=X ^{2011}+X ^{1783}-3X ^{1707}+2X ^{341}-3X ^{2}-3\]. Find the remainder, without using a calculator, when you divide P(x) by \[X ^{3}-X\].

OpenStudy (anonymous):

Okay have you divided polynomials before?

OpenStudy (anonymous):

If you like we can go through this step by step.

OpenStudy (anonymous):

Hmmm... don't think so :)

OpenStudy (anonymous):

yes, that would be helpful!!!

OpenStudy (anonymous):

Okay the first step is to take the X^3 and find out how many times it goes into x^{2011}. _____________________________________________________ X^3−X |X^{2011}+X^{1783}−3X^{1707}+2X^{341}−3X^2−3 Do you know what the answer is?

OpenStudy (anonymous):

\[x^mx^n=x^{m+n}\]

OpenStudy (anonymous):

not sure what the answer is; honestly...

OpenStudy (anonymous):

that's okay, \[x^3x^n=x^{2011}\]\[x^n=x^{2011-3}=x^{2008}\] So it is x^{2008} X^{2008} _____________________________________________________ X^3−X |X^{2011}+X^{1783}−3X^{1707}+2X^{341}−3X^2−3 Now that we know this we need to times x^{2008} by x^3-x, can you do this step?

OpenStudy (anonymous):

if \[x ^{3}-x ^{n}=2008\] then \[x ^{n}=2005\]?

OpenStudy (anonymous):

not x3-xn, but x3xn... sorry

OpenStudy (anonymous):

\[x ^{3}x ^{n}=2008\]

OpenStudy (anonymous):

yes that working is correct but we need to multiply \[x^{2008}(x^3-x)\]

OpenStudy (anonymous):

I'm not sure if you've done long division before but we are using this method. it's the easiest way to do these polynomial questions :)

OpenStudy (anonymous):

oh... ok

OpenStudy (anonymous):

would it be better if I did another example first and then continue with this one so you know the method?

OpenStudy (anonymous):

yeah, that would would be cool... thanks

OpenStudy (anonymous):

okay let's do this one __________________ X^2+3 |X^3-5X^2+3X-15

OpenStudy (anonymous):

Step 1 : find out how many x^2 go into x^3 X^{3-2}=X^1 or X X __________________ X^2+3 |X^3-5X^2+3X-15 Step 2: multiply x by X^2+3, put this underneath the division X(X^2+3)=X^3+3X X __________________ X^2+3 |X^3-5X^2+3X-15 X^3+3X Step 3: minus the two under the division sign X __________________ X^2+3 |X^3-5X^2+3X-15 -( X^3 +3X) ---------------------- 0 - 5X^2 +0 -15 -5X^2-15

OpenStudy (anonymous):

understand so far... thanks :)

OpenStudy (anonymous):

Step 4: Redo step 1-3 but for -5X^2 -5X^{2-2}=-5X^0 or -5 X -5 __________________ X^2+3 |X^3-5X^2+3X-15 -5(X^2+3)=-5X^2-15 X - 5 __________________ X^2+3 |X^3-5X^2+3X-15 -( X^3 + 3X) ---------------------- 0 - 5X^2 +0 -15 -5X^2-15 -5X^2-15 ------------------------ 0+0 So therefore X^2+3 goes into X^3-5X^2+3X-15, X - 5 times with no remainder.

OpenStudy (anonymous):

Would you like to try your question again first and I'll see if it's right or if you're on the right track.

OpenStudy (anonymous):

yeah, i'll do that :)

OpenStudy (anonymous):

\[x ^{2011}\]\[x ^{2011}+x ^{1783}-3x ^{1707}+3x ^{2}-3\]

OpenStudy (anonymous):

\[x ^{2011-3}=x ^{2008}\]

OpenStudy (anonymous):

Yes that's good but I think we will have to try another method to solve this because you will have to do long division by hand way too many times. Just give me a minute to think about it

OpenStudy (anonymous):

What we are going to do is separate this big question into smaller ones and do the long division on those.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{X^{2011}}{X^3-X}+\frac{X^{1783}}{X^3-X}- \frac{3X^{1707}}{X^3-X}+\frac{2X^{341}}{X^3-X}-\frac{3X^2}{X^3-X}-\frac{3}{X^3-X}\] So are problem is now like this, much easier to handle. We are doing the exact same method of solving just smaller scale. try the first fraction

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hold on I'm going to see if anyone has an idea on how to do this one. I thought separating it would help but we have the same problem. Sorry

OpenStudy (anonymous):